OpenStudy (bekkah323):

Rationalize the denominator. Assume that all expressions under radicals represent positive numbers.

OpenStudy (bekkah323):

\[\sqrt[3]{\frac{ 4 }{ 5x^5y^2 }}\]

OpenStudy (anonymous):

multiply top and bottom by \(\sqrt[3]{25xy}\)

OpenStudy (bekkah323):

where did you get that?

OpenStudy (anonymous):

because \(5\times 25=5^3\) and \(\sqrt[3]{5^3}=5\)

OpenStudy (anonymous):

similarly \(x^5\times x=x^6\) and \(\sqrt[3]{x^6}=x^2\)

OpenStudy (anonymous):

and also \(y^2\times y=y^3\) and \(\sqrt[3]{y^3}=y\) so there will be no radical left in the denominator when you do this, just \[5x^2y\]

OpenStudy (bekkah323):

Im sorry, i'm still a little confused on how you got that.

OpenStudy (anonymous):

ok lets go slow your have to get rid of the cubed root in the denominator right? i mean that is what "rationalize" means in this case

OpenStudy (bekkah323):

ok, that part makes sense!

OpenStudy (anonymous):

so you do this by multiplying the denominator by whatever will make the inside stuff in to perfect cubes

OpenStudy (bekkah323):

what do you mean perfect cubes?

OpenStudy (anonymous):

\(y^3\) is a perfect cube it is the cube of \(y\)

OpenStudy (anonymous):

\(125\) is a perfect cube, it is the cube of \(5\)

OpenStudy (bekkah323):

ahhh ok

OpenStudy (anonymous):

and finally \(x^6\) is also a perfect cube, it is the cube of \(x^2\)

OpenStudy (anonymous):

now if you look above at the instructions i wrote, i hope the reasoning is more clear if not, let me know